OpenStudy (missy1999):

PLEASE HELP! Medal + Fan! What is the equation in point-slope form of the line passing through (−1, 3) and (1, 7)? a. y − 7 = 4(x − 1) b. y − 7 = 2(x − 1) c. y − 3 = 2(x − 1) d. y − 3 = 4(x + 1)

OpenStudy (eskelle):

do you know what point slope form is?

OpenStudy (missy1999):

I got y - 3 = 2(x + 1) but that's not an answer. and when I graphed it, y − 7 = 2(x − 1) passed through the points (−1, 3) and (1, 7)

OpenStudy (missy1999):

yes, it's y - y1 = m(x - x1)

OpenStudy (eskelle):

so you can plug in the information you know

OpenStudy (missy1999):

I feel like the answers to the question may be messed up, because (1, 7) are x2 and y2, not x1 and y1, so they wouldn't be plugged into the equation. so if the answer is B, you'd have to be used x2 and y2, not x1 and y1 also, it's y - 3 = 2(x + 1) plugged in

OpenStudy (missy1999):

but that's not an option

OpenStudy (eskelle):

hmm. i'm not sure. maybe contact your teacher

OpenStudy (missy1999):

okay I will, thank you

OpenStudy (mathmate):

@missy1999 ONE of the answers is correct. Your equation is also correct, but apparently does not correspond to any of the answer choices.

OpenStudy (mathmate):

The secret is that if you expand the correct answer, and yours, they will be identical. The reason for the difference is because in the point slope form, you can have the line pass through any one of the two given points. You and the correct answer apparently chose different points, so they \(appear\) to be different. That's why if you expand the equations, they will be identical. What to do? Use the same slope (that you calculated) and pass through the other point, then you will have an equation identical to one of the choices.